hartnn (hartnn):

Maybe a Math Question. If it took me 1 week to reach 60 score and another 3 weeks to reach 75 score, then how many more weeks are required to reach the score of 1.90 2.99 3.100 I know its exponential function but i don't know exact function.Plz Help!!

OpenStudy (anonymous):

Actually, it depends on u..........lol

hartnn (hartnn):

if i go at same rate??

OpenStudy (lgbasallote):

3. 3 years

OpenStudy (anonymous):

Let y=ax^b

OpenStudy (anonymous):

Now, find a and b

OpenStudy (anonymous):

1. 2 weeks 2. 3 weeks 3. i dont know

hartnn (hartnn):

how can u be so sure?@lgbasallote and no direct answers!! please explain. read code of conduct.

OpenStudy (lgbasallote):

lol

OpenStudy (lgbasallote):

i know because in 3 years im quitting OS

hartnn (hartnn):

wow , 99 in 3 weeks!! thnx @mukushla

OpenStudy (anonymous):

Oh.... since u said at the same rate........... y=mx+c

OpenStudy (anonymous):

find m and c

hartnn (hartnn):

but exponential function it is,not linear.

hartnn (hartnn):

maybe y=ax^b+c.........

OpenStudy (anonymous):

But u said at the same rate and in exponential function rate is not constane

hartnn (hartnn):

then i am willing to go at exponential rate....whatever it takes....

OpenStudy (arlen):

for 90 it'll take exactly four more weeks regardless of number of medals earned

OpenStudy (lgbasallote):

\[x = x_0 e^{kt}\] \[\large 75 = 60e^{k (3-1)}\] so \[ k = 0.1116\] \[90 = 60e^{0.1116t}\] so \[t = 3.63 weeks\] so... .63 more weeks and you'll be 90

OpenStudy (lgbasallote):

0.63 more weeks *

hartnn (hartnn):

only 0.63 !! not even 1 week more....awesome!

hartnn (hartnn):

@arlen how??

OpenStudy (lgbasallote):

sadly SS is not about you..it's about others lol

OpenStudy (anonymous):

From 60 it took 3 weeks or from 0 it took 3 months to reach 75

OpenStudy (anonymous):

I mean 3 weeks

hartnn (hartnn):

from 60 it took me 3 weeks,not from 0

OpenStudy (arlen):

well this has been programmed keeping your lust of being a guru in mind...and to be a guru or having score of 99...you will have to keep logging in ...and if time spent on this is more, surely you'll be rewarded....cuz they have got to show maximum nuber of people online

OpenStudy (arlen):

NVM :)

hartnn (hartnn):

i got what u said.

OpenStudy (arlen):

i mean lust of everyone of us

hartnn (hartnn):

but frankly, i don't have any lust in becoming guru.

OpenStudy (arlen):

oh never mind...that was a bad joke

hartnn (hartnn):

i come here coz i love to solve math questions....medals are just toppings on ice-cream.

OpenStudy (anonymous):

According to my calculation Around 3.5 weeks to reach 90 AND another 3 weeks to reach 99.......... AND AS @mukushla I DONT KNOW WHEN WILL U REACH 100..... IT may take years

hartnn (hartnn):

lets see whose prediction is most accurate..........

OpenStudy (anonymous):

someone give a medal to saura...

hartnn (hartnn):

done.

OpenStudy (anonymous):

we are all 1 in this question...

hartnn (hartnn):

yeah,that looks nicer....

OpenStudy (anonymous):

THANX @mukushla and @hartnn

hartnn (hartnn):

ok, lets get back to work.....of helping ppl......

OpenStudy (lgbasallote):

it will take 5 weeks for @hartnn to reach 100 lol

hartnn (hartnn):

soooo early!!!!

OpenStudy (anonymous):

lol

OpenStudy (lgbasallote):

hartnn is good

hartnn (hartnn):

good in the sense ??

hartnn (hartnn):

took me 12 days to reach 90.............short lived golden days. green is my fav. color! and i m green now :D @mukushla 's calculation was most accurate, 2 weeks.